MODS1 Dark Commissioning and Spring SV Run: UTC 2011 June 1

Observer(s): R. Pogge (OSU), M. Wagner (OSU/LBTO), O. Kuhn (LBTO)
Telescope Operator: D. Gonzalez-Huerta (LBTO)
Telescope Support: J. Little (LBTO)
Instrument Support: none

Summary

Preparations

Instrument is as we left it last night. LN2 dewars filled a little after noon today, and this will be the last filling before it comes off the telescope. Mark Wagner and I will put on optics covers and disconnect MODS1 from LDG services tomorrow to prep it for removal in 3 weeks.

Details

All times are UTC unless otherwise noted.

0130 - Heavy high overcast rolled in late afternoon and it is still over us. Humidity is low (24%) but maybe some verga in the distance so we are not going to open right now. Winds are out of the south 5-10m/s. Warmer than yesterday, currently 12°C

0400 - A bit of clearing, but still clouds threatening on the satellite. David has opened up and we're equilibrating.

0405 - Pointing Setup
   IE = -58
   CA = +68

0410 - BS9152 Collimation Star. Seeing about 1-arcsec, quick convergence

J1420-1620 - science target

0420 - J1420-1620 (SV_DeRosa)

Acquired - switching to LS5x60x1.2 slit because seeing is ~1.1-1.2 arcsec. Some high clouds rolling through, so transparency leaves much to be desired and will be variable.

First acquisition image at 1423 (long preset), switched in 1.2 for 0.8 slit. Target nearly exactly on slit, moved +4arcsec in Y to center. Started integrations at 0428. Time on target is 2h, in 6x1200s chunks. This gives us an edge if we have to abort, and 1200s is about as long as you want in red due to an increased CR rate with this thick CCD.

0453 - first spectra hit the disk, can easily see Ly-alpha at the QSO redshift (z=5.73) at around 8100A. Nice. We'll be here for a while.

0536 - first hour of two is done. Median of 3 looks very good - tracking and flexure compensation are spot on. Seeing has been 1.1 arcsec, and we've had clouds going through, integrated guide-star counts have fluctuated quite a bit.

0545 - seeing blowing up about midway through exposure 4, up to 2-arcsec and variable. Ick. Sounds like the wind is coming up, and we're facing into the wind. Oh well.

0557 - Seeing has gone >2-arcsec and wind is gusting. We are aborting the sequence at the end of exposure 4. First hour was very nice and should give the PIs a taste of things, modulo accounting for the fact we were looking through clouds.

0605 - David swung the LBT around to get a better wind reading (we were face into the wind before). Seeing sustained 15m/s with gusts to 18m/s. Not at the limits, but much stronger than before. That's why the seeing went to hell.

J1717+2814 - science target

0620 - J1717+2814 SV_Howk. LS5x60x1.0, Acquired with back to the wind, seeing better here, about 1-arcsec, so we can go for it.

0624 - finished preset and acquisition, on starting the observation the wind gusted over limits and David shut us down. Bummer.

During shutdown we got alarms on the left shutter door. David et al. are investigating...

0800 - Latch pin problem solved. Clouds moving back in, thicker this time, wind is at ~16m/s sustained, but no big gusts. No cause to open yet...

-- RichardPogge - 01 Jun 2011
Topic revision: r2 - 01 Jun 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback